OpenStudy (anonymous):

find all real numbers x that x^4+6x^2-72=0

Directrix (directrix):

±√6

hero (hero):

If x^2 = y then y^2+6y-72=0 so y^2+12y-6y-72=0 thus y(y+12)-6(x+12)=0 therefore (y+12)(y-6)=0 since y+12 = 0 yields negative values of y: only y - 6 = 0 is true replacing y = x^2 x^2 - 6 = 0 x^2 = 6 x = + or - sqrt(6)